OpenStudy (anonymous):

What is the surface area of a rectangular prism that has these dimensions? 2 cm x 3 cm x 5 cm

OpenStudy (anonymous):

2*[(2 * 3) + (3 * 5) + (5 * 2)] = 2*[6 + 15 + 10] = 2*31 = 62 sq cm

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome ;-)

OpenStudy (anonymous):

???? what ???